OpenStudy (anonymous):

Help Partial Fractions \[\int\limits_{}^{}\frac{ x^2 }{ (x-1)(x^2+4x+5) }dx\]

OpenStudy (anonymous):

I have it all the way down to \[\int\limits_{}^{}\frac{ \frac{ 1 }{ 10 } }{ (x-1) }+\frac{ \frac{ 9 }{ 10 }+\frac{ 1 }{ 2 } }{ (x ^{2}+4x+5) }\]

OpenStudy (anonymous):

Do you want the heavyside method or the other one. I forgot the name sorry

OpenStudy (anonymous):

I'm really just having trouble simplifying from there to get to the answer

OpenStudy (anonymous):

and there should be an x after the 9/10

OpenStudy (anonymous):

Ok give me a sec I'll see if what u have is right

OpenStudy (anonymous):

sorry i'm forgetting my integration skills i might come back to this later

OpenStudy (anonymous):

np I've been there